OpenStudy (anonymous):

7x + 2y = 41 8x + 8y = 64

OpenStudy (texaschic101):

8x + 8y = 64 -- divide both sides by 8 x + y = 8 ---> x = 8 - y now sub 8 - y in for x in the other equation 7x + 2y = 41 7(8 - y) + 2y = 41 -- distribute through the parenthesis 56 - 7y + 2y = 41 -- subtract 56 from both sides -7y + 2y = 41 - 56 -- combine like terms -5y = -15 -- divide both sides by -5 y = 3 now sub 3 in for y in either equation to find x x + y = 8 x + 3 = 8 x = 8 - 3 x = 5 check.. 8x + 8y = 64 8(5) + 8(3) = 64 40 + 24 = 64 64 = 64 (correct) x = 5 and y = 3